# A quasi-optimal upper bound for induced paths in sparse graphs

### v1

@@ Abstract
In 2012, Nešetřil and Ossona de Mendez proved that graphs of bounded degeneracy that have a path of order $n$ also have an induced path of order $\Omega(\log\log n)$. In this paper we give an almost matching upper bound by describing, for arbitrarily large values of $n$, a 2-degenerate graph that has a path of order $n$ and where all induced paths have order $O(\log\log n\cdot\log\log\log n)$.
@@ 1 Introduction
In this work, we aim to understand the relation between long paths and induced paths in graphs. Let us denote by $\operatorname{\sf lp}(G)$ and $\operatorname{\sf lip}(G)$ the maximum orders of a path and of an induced path in a graph $G$. We trivially always have $\operatorname{\sf lip}(G)\leq\operatorname{\sf lp}(G)$ since an induced path is in particular a path. Conversely, in which graph classes is $\operatorname{\sf lip}$ lower-bounded by (a function of) $\operatorname{\sf lp}$? More formally, we are interested in graph classes $\mathcal{G}$ for which there is an unbounded function $f$ satisfying, for every $G\in\mathcal{G}$, $\operatorname{\sf lip}(G)\geq f(\operatorname{\sf lp}(G))$. This topic is studied since the 80’s at least when Galvin, Rival, and Sands proved the following general result.
Theorem 1.1 ([GRS82]).
@@
For every $t$ there is an unbounded function $f$ such that if $G$ is a $K_{t,t}$-subgraph free graph that has a path of order $n$, then $G$ has an induced path of order at least $f(n)$.
@@
Actually, the authors of the above result were not interested in an explicit expression of the function $f$, but it can be extracted from their proof and the currently best-known bounds on the multicolor Ramsey numbers that $f(n)=\Omega((\log\log\log n)^{1/3})$. This result has been improved to $f(n)=\Omega((\log\log n)^{1/5})$ in [DER24]. Very recently, Hunter, Milojević, Sudakov, and Tomon gave an elegant and short proof that $f(n)=\Omega(\log\log n/\log\log\log n)$ [HMST24]. At the time of writing this is the best-known lower-bound on $f$ and no matching upper bound is known.
On the other hand, there are graph classes where a statement such as Theorem 1.1 does not hold. This is the case of any class containing arbitrarily large cliques or balanced bicliques. Indeed, such graphs have a path containing all their vertices, yet no induced path of order more than 2 or 3, respectively. Note that an hereditary graph class that does not contain arbitrarily large cliques or balanced bicliques is $K_{t,t}$-subgraph free for some $t$. Hence, for hereditary graph classes Theorem 1.1 is best possible in the sense that the $K_{t,t}$-subgraph freeness assumption cannot be dropped.
@@
Let us say that a graph $G$ is $k$-degenerate if every subgraph of $G$ has a vertex of degree at most $k$. Note that $k$-degenerate graphs are $K_{k+1,k+1}$-subgraph free, hence they are covered by the aforementioned results. In the Sparsity textbook [NOdM12], Nešetřil and Ossona de Mendez gave a simple yet clever proof of the following. Unless otherwise specified, all logarithms are binary.
Theorem 1.2 ([NOdM12]).
@@
The authors of the above explicitly asked the question of finding the optimal bound.
Question 1.3 ([NOdM12, Problem 6.1]).
@@
We note that several recent papers also study the length of induced paths in graph classes such as planar graphs or graphs of bounded treewidth, see [DER24] for an overview.
Since [NOdM12] appeared, no improvement over the bound of Theorem 1.2 was made. On the other hand, in 2023 the second and the third author gave the following upper bound: there are 2-degenerate graphs that have a path of order $n$ and where all induced paths have order $O((\log\log n)^{2})$ [DR24]. Observe that 2-degenerate graphs are also $K_{t,t}$-subgraph free for every $t\geq 3$, hence this also provides an upper bound on the function of Theorem 1.1.
Contribution and techniques.
@@
This is an improvement over the $O((\log\log n)^{2})$ bound of [DR24] and answers Question 1.3 up to a $\log\log\log n$ factor. Note that 2-degenerate graphs are also $k$-degenerate for any $k\geq 3$ hence the above result is indeed an upper bound on the funtion of Theorem 1.2, for any value of $k$. As a consequence, the bounds on the functions of the aforementioned theorems become:
  • for the function $f$ of Theorem 1.1, there are two constants $c_{1},c_{2}$ such that for every integer $n$,
    $c_{1}\cdot\frac{\log\log n}{\log\log\log n}\leq f(n)\leq c_{2}\cdot\log\log n\cdot\log\log\log n;$
  • for the function $f_{k}$ of Question 1.3, there is a constant $c$ such that for every integer $n$,
    $\frac{\log\log n}{\log(k+1)}\leq f_{k}(n)\leq c\cdot\log\log n\cdot\log\log\log n.$
@@
The first step is the explicit construction of the graph whose existence is posited by Theorem 1.4. This graph should be 2-degenerate, contain a long path $P$ and no long induced paths. Observe that vertices outside of $P$ are not useful to show any of the desired properties. So actually our construction will not contain such vertices: the path $P$ will span all the vertices, i.e., it will be a Hamiltonian path. More precisely, for every integer $\ell$, we construct a 222-degenerate graph $G$ of order $\smash{\Omega\big{(}2^{2^{\ell}}\big{)}}$ that admits a Hamiltonian path and in which all induced paths have order $O(\ell\log\ell)$. The different steps of the construction can be sketched as follows (see Figure 1 for an overview):
@@
    We start with a complete binary tree on $\smash{2^{2^{\ell}-1}}-1$ nodes, and that we call the skeleton-tree. This tree is obviously 222-degenerate, has the desired number of vertices, but there are two drawbacks: it has no Hamiltonian path and its longest induced paths too long (or order $\Theta(2^{\ell}))$.
@@
    This skeleton-tree is transformed into a a graph called ribbed-tree obtained by adding a number of barriers. Informally, barriers are paths replacing the edges of the skeleton-tree together with a set of edges called ribs that connect nodes of such paths to certain well-chosen ancestor nodes in the tree. The role of these barriers is to block induced paths and they do so without increasing the degeneracy. Their definition uses an auxiliary tree called index-tree and certain walks on this tree. While it can be shown that the induced paths have now order $O(\ell\log\ell)$ as desired (for an appropriate choice of the index-tree), the obtained graph does not have a Hamiltonian path.
@@
    Finally, we apply on the ribbed-tree an operation called blow-up that results in a graph $G$ that contains a Hamiltonian path, remains 222-degenerate, has $\smash{\Omega(2^{2^{\ell}})}$ vertices (hence a path of the same order), and where induced paths have order ${O(\ell\log\ell)}$, as desired.
@@
The rest of the paper is organized as follows. In Section 2 we introduce the standard notations that we use in this paper. In Section 3 we detail the construction of our graph following the different steps highlighted in Figure 1. At the end of this section is argued that the obtained graph contains a Hamiltonian path, and that it is 222-degenerate. In Section 4 we characterize the shape of particular induced paths in this graph, and show that their order is function of the barrier’s size. In Section 5 we upper bound the barrier’s size, and prove Theorem 1.4 in Section 6.
@@ 2 Preliminaries
We assume the reader familiar with standard graph theory terminology, and only recall the different types of trees we consider in this paper, which are defined as follows.
@@
[…]
@@ 3 The construction
We describe the construction of a family of graphs satisfying Theorem 1.4. Namely, for each integer $\ell$, we show how to construct a 222-degenerate graph $G$ with $\smash{\Omega(2^{2^{\ell}})}$ vertices, containing a Hamiltonian path, and no induced path of order more than $c\cdot\ell\log\ell$ for some integer $c$.
The construction of $G$ consists of different steps which are split into several sections, each involving an auxiliary structure. First, an underlying tree structure called skeleton-tree is described in Section 3.1. This tree is extended into a ribbed-tree by integrating a number of “barriers” aimed at blocking induced paths in Section 3.3. This later notion of barrier involves trees and intricate sequences of integers that are described before, in Section 3.2. Finally, the obtained graph is modified to make it traceable and 222-degenerate in Section 3.4.
@@ 3.1 Skeleton-trees
  • The tree ${\sf S}_{1}$ is the single-vertex graph, and its root is its unique vertex. The rank of this vertex is 111 and its zone is ${\sf S}_{1}$ itself.
@@
      We start from a node $r$, which is the root of ${\sf S}_{\ell}$, two copies of ${\sf S}_{\ell-1}$ of respective roots $r_{1}$ and $r_{2}$, and make $r$ adjacent to $r_{1}$, its left child, and to $r_{2}$, its right child.
@@
      Then, for every leaf $t$ of the obtained tree, we create two additional copies of ${\sf S}_{\ell-1}$ of respective roots $r^{\prime}_{1}$ and $r^{\prime}_{2}$, and make $t$ adjacent to $r^{\prime}_{1}$, its left child, and to $r^{\prime}_{2}$, its right child.
    This completes the construction of ${\sf S}_{\ell}$. The rank of $r$ is $\ell$, its zone is ${\sf S}_{\ell}$, and the ranks and zones of the other nodes are those that were assigned to them in ${\sf S}_{\ell-1}$.
For a node $t$ of ${\sf S}_{\ell}$, the function ${\sf zone}\text{-}{\sf ancestor}_{t}$margin: ${\sf zone}\text{-}{\sf ancestor}_{t}$ maps any integer $i$ with ${\sf rank}(t)<i\leq\ell$ to the unique node $s$ such that $t\in{\sf zone}(s)$ and ${\sf rank}(s)=i$. Moreover, we define margin: ${\sf zone}_{t}$ ${\sf zone}_{t}(i):={\sf zone}({\sf zone}\text{-}{\sf ancestor}_{t}(i))$ if $i\leq\ell$. See Figure 2 for an illustration of this terminology.
@@
Adjacent nodes in ${\sf S}_{\ell}$ are either of consecutive rank, or one of the two has rank 111, a point that will be important in the remaining of the proof.
@@
By construction, $0pt({\sf S}_{1})=|V({\sf S}_{1})|=1$. For any $\ell\geq 2$,
$0pt({\sf S}_{\ell})=2\cdot 0pt({\sf S}_{\ell-1})+1$
@@ 3.2 Index-trees and index-barriers
Let us fix an $\ell$-index-tree ${\sf T}$. Given an index $i$ of ${\sf T}$ we denote by $i^{-}$ its left child, and by $i^{+}$ its right child;margin: $i^{-}$, $i^{+}$ the subtree rooted at $i$ is denoted by ${\sf T}(i)$,margin: ${\sf T}(i)$ and by ${\sf T}(i^{-})$ and ${\sf T}(i^{+})$ are thus meant the subtrees rooted at $i^{-}$ and $i^{+}$, respectively. If the left or right child does not exist these notations refer to the empty set and the empty graph.
@@
For any $1\leq i\leq\ell$, the full index-barriermargin: full index-barrier of $i$ denoted $B(i)$ is the word over the alphabet $\{1,\dots,i\}$ defined by:margin: $B(i)$
@@
This alternative but equivalent view will be of interest later, in particular to make the following basic observations as well as to prove a number of properties on factors of $B(\ell)$.
@@
If two indices $i,j\in\left\{1,\dots,\ell\right\}$ appear in some factor $W$ of $B(\ell)$, then every index in the shortest $i$$j$ path in ${\sf T}$ appears in $W$ as well.
@@
Let $k,i,j\in\left\{1,\dots,\ell\right\}$ be such that $k>i>j$ and $i,j\in B(k)$. Then the first occurrence of $i$ appears before the first occurrence of $j$ in $B(k)$.
@@
Let $i,b\in\{1,\dots,\ell\}$. If $i<b$ then either $i$ lies in ${\sf T}(b^{-})$, or it lies in ${\sf T}(k^{+})$ for some ancestor $k$ of $b$ with possibly $k=b$.
@@
Consider a factor $W$ of $B(\ell)$ from an occurrence of $i^{\prime}$ to an occurrence of $j$. Without loss of generality we may assume that none of $i^{\prime}$ and $j$ appears as an internal letter of $W$ (otherwise we may consider the shorter factor using that letter). Then by definition of a full index-barrier, $W$ is a prefix of $i^{\prime}\circ B(i)\circ i^{\prime}$. Applying Observation 3.6 with $i^{\prime}$ instead of $k$, and every $j^{\prime}\in\left\{i,\dots,j+1\right\}$ instead of $i$, we obtain that all integers between $i$ and $j+1$ appear before the first occurrence of $j$, hence in $W$, as claimed. ∎
@@
Note that the index-barrier $B(i,j)$ is defined for $i>j$ only. It contains one occurrence of $j$ and possibly several occurrences of $i$. Moreover, by Observation 3.6, the occurrence of $j$ in $B(i,j)$ is in fact the first occurrence of $j$ in $B(\ell)$. In the remaining of the paper, two particular types of index-barriers will be considered, namely, those of the form
  • $B(i,j)$ for $i=j+1$; and
  • $B(i,1)$ for $i\geq 2$.
@@
For any $i,j\in\left\{1,\dots,\ell\right\}$ with $i>j$, $B(i,j)$ contains all indices from $i$ to $j$, no internal occurrence of $j$, and it does not contain indices smaller than $j$. In particular, any prefix of $B(i,1)$ from $i$ to the first occurrence of $j<i$ satisfies these properties.
@@
For any $i\in\left\{2,\dots,\ell\right\}$, $B(i,i-1)$ contains indices greater or equal to $i-1$, and it does not contain any internal occurrence of $i$ nor $i-1$.
@@
The fact that $B(i,i-1)$ does not contain internal occurrences of $i-1$ follows by definition, and the fact that it only contains indices greater than $i-1$ follows by Observation 3.11. Let us show that no internal occurrence of $i$ appears in $B(i,i-1)$.
@@
Suppose that $i$ is not a leaf. All occurrences of $i$ appears in a factor $B(i)$ by definition. The first occurrence of $i$ is therefore in $B(i)$. If $i$ has two children, $i-1$ is its left child, and $B(i)=i\circ B(i-1)\circ i\circ B(i^{+})\circ i\circ B(i-1)\circ i$ and $B(i-1)$ start with $i-1$. Therefore $B(i,i-1)=i\circ i-1$. Similarly, if $i$ has one child, $B(i)=i\circ B(i-1)\circ i$ and therefore $B(i,i-1)=i\circ i-1$.
Suppose that $i$ is a leaf. To prove this case, we first need to observe that if an index $i^{\prime}$ is the smallest that appears in $B(j)$, then it appears exactly once. This can be shown by the following induction. It is true for $j$ a leaf as in this case $B(j)=j=i^{\prime}$. If $j$ has two children then $i^{\prime}$ is in ${\sf T}(j^{-})$ and it is the smallest of $T(j^{-})$ therefore it appears once in $B(j^{-})$ by induction, and so once in $B(j)=j\circ B(j^{-})\circ j\circ B(j^{+})\circ j\circ B(j^{-})\circ j$. Similarly if $j$ have one child $B(j)=j\circ B(j-1)\circ j$ and the smallest index appears once.
Now let $j$ be the smallest index such that $i$ and $i-1$ belong to ${\sf T}(j)$. Then $j$ has two children. Since $i$ is a leaf, $i\in{\sf T}(j^{-})$ and $i-1$ is the right child of $j$. Since $i$ is the smallest index of ${\sf T}(j^{-})$ it appears once in $B(j^{-})$. So $B(j)=j\circ B(j^{-})\circ j\circ B(i-1)\circ j\circ B(j^{-})\circ j$. Since $B(i,i-1)$ is a factor of $B(j)$ and since $B(j^{-})$ contains only one occurrence of $i$, so does $B(i,i+1)$. ∎
@@
Let $i,b\in\left\{1,\dots,\ell\right\}$ be such that $i<b$. Then all the factors of $B(\ell)$ starting with some $b$, ending with some $b$, containing at least one occurrence of $i$, and no other occurrence of $b$, are equal.

### v2

@@ Abstract
In 2012, Nešetřil and Ossona de Mendez proved that graphs of bounded degeneracy that have a path of order $n$ also have an induced path of order $\Omega(\log\log n)$. In this paper we give an almost matching upper bound by describing, for arbitrarily large values of $n$, 2-degenerate graphs that have a path of order $n$ and where the longest induced paths have order $O((\log\log n)^{1+o(1)})$.
@@ 1 Introduction
Ramsey’s Theorem, proved in 1930, is a cornerstone of extremal combinatorics. In in its finite form, it states that every large enough graph excluding a given clique contains a large independent set. There are several analogues of this result where the desired objects are not independent sets but other induced substructures and where the considered graphs should satisfy some additional properties. A simple example is the following theorem for connected graphs.
Theorem 1.1 (see [DOO+96, (5.3)]).
@@
For every integer $t\geq 1$, there is an increasing function $f$ such that every connected $K_{t}$-free graph on $n$ vertices contains as induced subgraph $K_{1,f(n)}$ or $P_{f(n)}$.
There are also further extensions for other substructures than induced subgraphs and various requirements of connectivity, see [ALL22] and the references therein. In the 80’s, Galvin, Rival, and Sands proved the following Ramsey-type result for traceable graphs, i.e., graphs that have a Hamiltonian path.
Theorem 1.2 ([GRS82, Theorem 4]).
For every integer $t\geq 1$, there is an unbounded function $h_{t}$ such that every $K_{t,t}$-subgraph free traceable graph on $n$ vertices contains an induced path of order at least $h_{t}(n)$.
This can also be restated as follows: if a $K_{t,t}$-subgraph free graph has a path of order $n$, then it has an induced path of order at least $h_{t}(n)$. Note that cliques and bicliques do not satisfy such a statement: they have a path visiting all their vertices yet no induced path of order more than 2 or 3, respectively. Hence, for hereditary graph classes, being $K_{t,t}$-subgraph free is necessary in the above statement. [^1] If a graph contains a large biclique as a subgraph, it contains either a large biclique as induced subgraph, or a large clique (by Ramsey’s theorem).
The question motivating this work is to find good bounds on the function $h_{t}$ of Theorem 1.2. In addition to being interesting on its own, this question was asked in [GRS82] already in the special case $t=2$. As additional motivation, let us describe two other research lines that would benefit from progress on this question. The first one is algorithmic. Theorem 1.2 was rediscovered thirty years later in [ALR12] (with worse bounds) and used to design a parameterized algorithm for the Biclique problem. Any improved bound on the growth rate of $h_{t}$ would result in an improved complexity bound. The second one is related to treedepth, an important structural graph parameter in the theory of sparse graphs. [^2] Because we will not use treedepth in this paper, we refrain from providing the definition and refer the interested reader to [NO12, Chapter 6] instead. It is well-known that a graph with no path of order $n$ has treedepth at most $n$ (see, e.g., [NO12, Proposition 6.1]). Using Theorem 1.2 we can deduce the following induced analogue: a $K_{t,t}$-subgraph free graph with no induced path of order $h_{t}(n)$ has treedepth at most $n$. Here again improving bounds on $h_{t}(n)$ would strengthen the statement.
@@
In [GRS82], the authors of Theorem 1.2 did not provide an explicit expression of the function $h_{t}$, but it can be extracted from their proof and the currently best-known bounds on the multicolor Ramsey numbers that $h_{t}(n)=\Omega_{t}((\log\log\log n)^{1/3})$. This result has recently been improved to $h_{t}(n)=\Omega_{t}((\log\log n)^{1/5})$ by Duron, Esperet, and the last author [DER24]. Very recently, Hunter, Milojević, Sudakov, and Tomon gave in [HMS+24] an elegant and short proof that
$h_{t}(n)=\Omega_{t}(\log\log n/\log\log\log n).$
At the time of writing this is the best-known lower-bound on $h_{t}$ and no matching upper-bound is known.
@@
A related question was asked in the Sparsity textbook [NO12]. Let us say that a graph $G$ is $k$-degenerate if every subgraph of $G$ has a vertex of degree at most $k$. Note that $k$-degenerate graphs are $K_{k+1,k+1}$-subgraph free, hence they are covered by the results above. In the aforementioned book, Nešetřil and Ossona de Mendez gave a simple yet clever proof of the following statement. (Unless otherwise specified, all logarithms are binary.)
Theorem 1.3 ([NO12]).
@@
The also explicitly asked the question of finding the optimal bound.
Question 1.4 ([NO12, Problem 6.1]).
@@
More generally, for a graph class $\mathcal{C}$ we can define $f_{\mathcal{C}}$ as the maximum function such that, for every graph $G\in\mathcal{C}$ and every $n$, if $G$ has a path of order $n$ then $G$ has an induced path of order $f_{\mathcal{C}}(n)$. Several recent papers study the growth rate of $f_{\mathcal{C}}$ for various choices of $\mathcal{C}$ such as planar graphs or graphs of bounded treewidth, see [DER24] for an overview.
Since [NO12] appeared, no improvement over the lower bound of Theorem 1.3 was made. On the other hand, in 2023 the second and the third author gave the following upper bound: there are 2-degenerate graphs that have a path of order $n$ and where all induced paths have order $O((\log\log n)^{2})$ [DR24]. Observe that 2-degenerate graphs are also $K_{t,t}$-subgraph free for every $t\geq 3$, hence this also provides an upper bound on the function of Theorem 1.2.
@@
This is an improvement over the $O((\log\log n)^{2})$ bound of [DR24] and answers Question 1.4 up to a $\log\log\log n$ factor. Note that 2-degenerate graphs are also $k$-degenerate for any $k\geq 3$ hence the above result is indeed an upper bound on the funtion of Theorem 1.3, for any value of $k$. As a consequence, the bounds on the functions of the aforementioned theorems become:
  • for every integer $t\geq 3$, there is a constant $c_{t}>0$ such that for every integer $n\geq 16$ the function $h_{t}$ of Theorem 1.2 satisfies
    $c_{t}\cdot\frac{\log\log n}{\log\log\log n}\leq h_{t}(n)\leq c\cdot\log\log n\cdot\log\log\log n;$
  • for every integers $k\geq 1$ and $n\geq 4$, the function $f_{k}$ of Question 1.4 satisfies:
    $\frac{\log\log n}{\log(k+1)}\leq f_{k}(n)\leq c\cdot\log\log n\cdot\log\log\log n,$
where $c$ is the constant from Theorem 1.5.
Our techniques.
@@
The first step is the explicit construction of the graph whose existence is posited by Theorem 1.5. This graph should be 2-degenerate, contain a long path $P$ and no long induced paths. Observe that vertices outside of $P$ are not useful to show any of the desired properties. So actually our construction will not contain such vertices: the path $P$ will span all the vertices, i.e., it will be a Hamiltonian path. More precisely, for every integer $\ell$, we construct a $2$-degenerate graph $G$ of order $\smash{\Omega\big(2^{2^{\ell}}\big)}$ that admits a Hamiltonian path and in which all induced paths have order $O(\ell\log\ell)$. The different steps of the construction can be sketched as follows (see Figure 1 for an overview):
@@
    We start with a complete binary tree on $\smash{2^{2^{\ell}-1}}-1$ nodes, and that we call the skeleton-tree. This tree is obviously $2$-degenerate, has the desired number of vertices, but there are two drawbacks: it has no Hamiltonian path and its longest induced paths are too long (of order $\Theta(2^{\ell}))$.
@@
    This skeleton-tree is transformed into a a graph called ribbed-tree obtained by adding a number of barriers. Informally, barriers are paths replacing the edges of the skeleton-tree together with a set of edges called ribs that connect nodes of such paths to certain well-chosen ancestor nodes in the tree. The role of these barriers is to block induced paths. They are defined in a way so that the obtained graph remains 2-degenerate. Their definition uses an auxiliary tree called index-tree and certain walks on this tree. While it can be shown that the induced paths have now order $O(\ell\log\ell)$ as desired (for an appropriate choice of the index-tree), the obtained graph does not have a Hamiltonian path.
@@
    Finally, we apply on the ribbed-tree an operation called blow-up that results in a graph $G$ that contains a Hamiltonian path, remains $2$-degenerate, has $\smash{\Omega(2^{2^{\ell}})}$ vertices (hence a path of the same order), and where induced paths have order $O(\ell\log\ell)$}, as desired.
@@
The rest of the paper is organized as follows. In Section 2 we introduce the standard notations that we use in this paper. In Section 3 we detail the construction of our graph following the different steps highlighted in Figure 1. At the end of this section is argued that the obtained graph contains a Hamiltonian path, and that it is $2$-degenerate. In Section 4 we characterize the shape of particular induced paths in this graph, and show that their order is function of the barrier’s size. In Section 5 we upper bound the barrier’s size, and prove Theorem 1.5 in Section 6.
@@ 2 Preliminaries
We assume that the reader is familiar with standard graph theory terminology, and only recall the different types of trees we consider in this paper, which are defined as follows.
@@
The depthmargin: depth of a node $s$ in a rooted tree is the order of the unique path from $s$ to the root, denoted $0pt(s)$; note in particular that the root has depth $1$. The depth of a rooted tree $T$, denoted $0pt(T)$, is the maximum order of a path from the root to the leaf.
@@ 3 The construction
We describe the construction of a family of graphs satisfying Theorem 1.5. Namely, for each integer $\ell$, we show how to construct a $2$-degenerate graph $G$ with $\smash{\Omega(2^{2^{\ell}})}$ vertices, containing a Hamiltonian path, and no induced path of order more than $c\cdot\ell\log\ell$ for some integer $c$.
The construction of $G$ consists of different steps which are split into several sections, each involving an auxiliary structure (see Figure 1). First, an underlying tree structure called skeleton-tree is described in Section 3.1. This tree is extended into a ribbed-tree by integrating a number of “barriers” aimed at blocking induced paths in Section 3.3. This later notion of barrier involves trees and intricate sequences of integers that are described before, in Section 3.2. Finally, the obtained graph is modified to make it traceable and $2$-degenerate in Section 3.4.
@@ 3.1 Skeleton-trees
  • The tree ${\sf S}_{1}$ is the single-vertex graph, and its root is its unique vertex. The rank of this vertex is $1$ and its zone is ${\sf S}_{1}$ itself.
@@
      We start from a node $r$, which will be the root of $\sf S$$_{\ell}$, two disjoint copies of ${\sf S}_{\ell-1}$ of respective roots $r_{1}$ and $r_{2}$, and make $r$ adjacent to $r_{1}$, its left child, and to $r_{2}$, its right child.
@@
      Then, for every leaf $t$ of the obtained tree, we create two additional disjoint copies of ${\sf S}_{\ell-1}$ of respective roots $r^{\prime}_{1}$ and $r^{\prime}_{2}$, and make $t$ adjacent to $r^{\prime}_{1}$, its left child, and to $r^{\prime}_{2}$, its right child.
    This completes the construction of $\sf S$$_{\ell}$. The rank of $r$ is $\ell$, its zone is $\sf S$$_{\ell}$, and the ranks and zones of the other nodes are those that were assigned to them in the copy of ${\sf S}_{\ell-1}$ they belong to.
For a node $t$ of $\sf S$$_{\ell}$, the function ${\sf zone}\text{-}{\sf ancestor}_{t}$margin: ${\sf zone}\text{-}{\sf ancestor}_{t}$ maps any integer $i$ with ${\sf rank}(t)<i\leq\ell$ to the unique ancestor $s$ of $t$ such that ${\sf rank}(s)=i$ and $t\in{\sf zone}(s)$. [^3] Note that in general $t$ might have several ancestors of rank $i$, but only one whose zone contains $t$. Moreover, we define margin: ${\sf zone}_{t}$ ${\sf zone}_{t}(i):={\sf zone}({\sf zone}\text{-}{\sf ancestor}_{t}(i))$ if $i\leq\ell$. See Figure 2 for an illustration of this terminology.
@@
The following simple observation will be important in the proof.
@@
Adjacent nodes in $\sf S$$_{\ell}$ are either of consecutive rank, or one of the two has rank $1$.
@@
By construction we have:
$\left\{\begin{array}[]{l}0pt({\sf S}_{1})=1\\ 0pt({\sf S}_{\ell})=2\cdot 0pt({\sf S}_{\ell-1})+1,\ \text{for any}\ \ell\geq 2.\end{array}\right$.
@@ 3.2 Index-trees and index-barriers
Let us fix an $\ell$-index-tree $\sf T$}. Given an index $i$ of $\sf T$} we denote by $i^{-}$ its left child, and by $i^{+}$ its right child;margin: $^{-}, i+i^{+}$ the subtree rooted at $i$ is denoted by ${\sf T}(i)$,margin: ${\sf T}(i)$ and by ${\sf T}(i^{-})$ and ${\sf T}(i^{+})$ are thus meant the subtrees rooted at $i^{-}$ and $i^{+}$, respectively. If the left or right child does not exist these notations refer to the empty set and the empty graph.
@@
For any index $i\in\left\{1,\dots,\ell\right\}$, the full index-barriermargin: full index-barrier of $i$, denoted $B(i)$, is the word over the alphabet $\{1,\dots,i\}$ defined by:margin: $B(i)$
@@
This alternative but equivalent point of view will be of interest later, in particular to make the following basic observations as well as to prove a number of properties on factors of $B(\ell)$.
@@
If two indices $i,j\in\left\{1,\dots,\ell\right\}$ appear in some factor $W$ of $B(\ell)$, then every index in the shortest iijj path in $\sf T$} appears in $W$ as well.
@@
Let $k,i,j\in\left\{1,\dots,\ell\right\}$ be such that $k>i>j$ and $i,j\in V({\sf T}(k))$. Then the first occurrence of $i$ appears before the first occurrence of $j$ in $B(k)$.
@@
Let $i,b\in\{1,\dots,\ell\}$. If $i<b$ then either $i$ lies in ${\sf T}(b^{-})$, or it lies in ${\sf T}(k^{+})$ for some ancestor $k$ of $b$ ($k=b$).
@@
Consider a factor $W$ of $B(\ell)$ from an occurrence of $i^{\prime}$ to an occurrence of $j$. Without loss of generality we may assume that none of $i^{\prime}$ and $j$ appears as an internal letter of $W$ (otherwise we may consider the shorter factor using that letter). Then by definition of a full index-barrier, $W$ is a prefix of $i^{\prime}\circ B(i)\circ i^{\prime}$. Applying Observation 3.6 with $i^{\prime}$ instead of $k$, and every $j^{\prime}\in\left\{i,\dots,j\right\}$ instead of $i$, we obtain that all integers between $i$ and $j+1$ appear before the first occurrence of $j$, hence in $W$, as claimed. ∎
@@
Note that the index-barrier $B(i,j)$ is defined for $i>j$ only. It contains one occurrence of $j$ and possibly several occurrences of $i$. Moreover, by Observation 3.6, the occurrence of $j$ in $B(i,j)$ is in fact the first occurrence of $j$ in $B(\ell)$. In the remaining of the paper, two particular types of index-barriers will be considered, namely, those of the form B(i1,i)B(i-1,i) and $B(i,1)$ for $i\geq 2$.
@@
From the definition and Proposition 3.8 we also have the following.
@@
For any $i,j\in\left\{1,\dots,\ell\right\}$ with $i>j$, $B(i,j)$ contains all indices from $i$ to $j$, no internal occurrence of $j$, and it does not contain indices smaller than $j$. In particular, any prefix of $B(i,1)$ ending at the first occurrence of $j$ satisfies these properties.
@@
For any $i\in\left\{2,\dots,\ell\right\}$, B(i,i1)B(i,i-1) only contains indices greater or equal to i1i-1, and it does not contain any internal occurrence of $i$ nor i1i-1.
@@
The fact that B(i,i1)B(i,i-1) does not contain internal occurrences of i1i-1 follows from the definition, and the fact that it only contains indices greater than i1i-1 is implied by Observation 3.11. Let us show that no internal occurrence of $i$ appears in B(i,i1)B(i,i-1).
@@
Suppose that $i$ is not a leaf. All occurrences of $i$ appears in a factor $B(i)$ by definition. The first occurrence of $i$ is therefore in $B(i)$. If $i$ has two children, i1i-1 is its left child, and $B(i)=i\circ B(i-1)\circ i\circ B(i^{+})\circ i\circ B(i-1)\circ i$ and B(i1)B(i-1) start with i1i-1. Therefore $B(i,i-1)=i\circ i-1$. Similarly, if $i$ has one child, $B(i)=i\circ B(i-1)\circ i$ and therefore $B(i,i-1)=i\circ i-1$.
Suppose that $i$ is a leaf. To deal with this case, we first need to prove the following statement: for every $j\in\left\{1,\dots,\ell\right\}$, if an index $i^{\prime}$ is the smallest that appears in $B(j)$, then it appears exactly once. This can be shown by the following induction on the depth of $j$. The statement is true when $j$ is leaf as in this case $B(j)=j=i^{\prime}$. If $j$ has two children then $i^{\prime}$ is in ${\sf T}(j^{+})$ (because the indices in ${\sf T}(j^{+})$ are smaller than those in ${\sf T}(j^{-})$). Since all indices of ${\sf T}(j^{+})$ appear in $B(j)$, we get that $i^{\prime}$ is the smallest index of $T(j^{+})$. Therefore, by induction, it appears once in $B(j^{+})$, and so only once in $B(j)=j\circ B(j^{-})\circ j\circ B(j^{+})\circ j\circ B(j^{-})\circ j$. Similarly, if $j$ has one child then by definition $B(j)=j\circ B(j-1)\circ j$ and by applying induction on j1j-1 we get that the smallest index appears only once.
Now let $j$ be the smallest index such that $i$ and i1i-1 belong to ${\sf T}(j)$ (that is, $j$ is the least common ancertor of $i$ and i1i-1). By construction of $\sf T$}, i1i-1 is not an ancestor of $i$ and since $i$ has no descendant, $j$ has two children. Since $i$ is a leaf, $i\in{\sf T}(j^{-})$ and i1i-1 is the right child of $j$. Since $i$ is the smallest index of ${\sf T}(j^{-})$ it appears once in $B(j^{-})$. So $B(j)=j\circ B(j^{-})\circ j\circ B(i-1)\circ j\circ B(j^{-})\circ j$. Since B(i,i1)B(i,i-1) is a factor of $B(j)$ and since $B(j^{-})$ contains only one occurrence of $i$, so does $B(i,i+1)$. ∎
@@
Let $i,b\in\left\{1,\dots,\ell\right\}$ be such that $i<b$. Then all the factors of $B(\ell)$ starting with $b$, ending with $b$, containing at least one occurrence of $i$, and no other occurrence of $b$, are equal.